A standing wave, generated on a string of length L, is represented by the following mathematical expression y(x,t) = 10 sin⁡(2πx) cos⁡(15πt). At what times would all the elements of the string have zero displacement [y(x,?) = 0]?


1
Expert's answer
2021-12-20T10:32:19-0500

"y(x,t) = 10 \\sin\u2061(2\u03c0x) \\cos\u2061(15\u03c0t)=0\\\\\\cos\u2061(15\u03c0t)=0\\\\15\u03c0t=\\frac{\\pi}{2}+\\pi n\\\\t=\\frac{1}{30}+\\frac{n}{15}"